### v1

# Inattention to States and Characteristics

Abstract
We study a generalization of the state-action rational inattention model with two dimensions of uncertainty: states and hedonic characteristics. The resulting conditional choice probability takes a familiar weighted multinomial logit form, but in contrast to the state-action model, is unique and is not a corner solution. The model imposes testable restrictions on the behavior of choice probabilities in markets with product entry. The optimal information policy can be found by optimizing subject to a marginal constraint, then finding the optimal marginal. The solution to the former is the Schrödinger bridge of an entropic optimal transport problem. The latter is solved by maximizing a strictly concave function (over a convex set) which envelops the objective.
Introduction
Consumer choice typically contains two dimensions of uncertainty: one concerns the product, and the second concerns utility-relevant future states of the world. When buying insurance, for example, the consumer faces uncertainty surrounding the contents of each insurance contract and their own future risk.
An increasingly popular framework for studying decision-making under uncertainty is the state-action rational inattention model [4]. In it, the decision-maker (DM) commits to an information policy consisting of a joint distribution $P$ over actions and states, subject to a Bayes plausibility constraint. The DM seeks to maximize expected utility less an information cost which is linear in the Shannon information between state and action. A relationship between the optimal conditional choice probability and the optimal marginal probability is known; however, the optimal marginal is not necessarily unique, and therefore neither is the optimal conditional.
We present a natural extension of the state-action model, which we call the state-characteristic model. In this model, the DM faces an exogenous prior distribution $\phi$ over characteristics $\xi$ – reflecting the inherent ubiquity of certain characteristics – and commits to a joint distribution $P$ over characteristics $\xi$ and states $\theta$. Mathematically, our model differs in only one respect: in addition to the usual Shannon information cost, we add a cost of diverging from $\phi$, interpretable as information gain from $P$ over $\phi$. This models, for instance, the fact that in a market full of lemons, one would need to acquire information to ensure that the likelihood of buying lemons is low.
The dual information costs account for the two dimensions of information gain that is observed by the econometrician when they observe the joint distribution over $(\xi,\theta)$. First, the econometrician observes there is information gain from $P(\xi|\theta)$ over $\phi(\xi)$ – that is, the choice distribution over characteristics is not what it would be if consumers chose at random. Second, the econometrician observes that the consumer’s choice of coverage level provides information about what their future losses will be – which differs from the prior $\mu$.
In section˜2, we motivate the need for the state-characteristic model by showing that the state-action model cannot account for both dimensions of information observed by the econometrician: modeling characteristics as actions fails to account for information gain from $P(\xi|\theta)$ over $\phi(\xi)$, while modeling characteristics as stochastic fails to account for information gain from $P(\theta|\xi)$ over $\mu(\theta)$. The state-characteristic model can be thought of as an interpolation between these two approaches which includes both information costs. We present the model in full generality in section˜3.
In our model, the relative costs for each dimension of information is parametrized by $\alpha\in[0,1]$. Our model thus nests the two aforementioned versions of the state-action model discussed in section˜2 at $\alpha=0$ and $\alpha=1$. As results have already been established for these cases, we focus on the case $\alpha\in(0,1)$.
Our main result, presented in section˜4, is very similar to the famous weighted multinomial logit (MNL) result [4]. Whereas their optimal conditional choice probability (CCP) satisfied
$P(a|\theta)=\frac{P(a)e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
for some normalizing partition function $Z$, our CCP satisfies
$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(a,\theta)/\lambda}}{Z(\theta;P)}$
One major advantage that our model ($\alpha\in(0,1)$) has is it always admits a unique solution whereas state-action ($\alpha=0$) does not. Therefore, even in cases where state-action seems to better model the situation, the state-characteristic model may nevertheless be useful for selecting from multiple solutions. Another distinction is that in state-action model, the DM forms (possibly non-unique) optimal “consideration sets” [1] – such that actions outside the set have $P(a)=0$ – whereas a key feature of the state-characteristic model is that no characteristic is completely ruled out.
This reflects the idea that consumers typically content themselves with being pretty sure of certain things; being truly 100% sure of anything requires an irrational amount of effort. For example, to acquire information as to how sausages are made, the conscious consumer may read the nutrition facts, the concerned consumer might read the ingredients, but nobody burglarizes a sausage factory with a spectrometer to make sure that the sausages are not produced in an environment with trace amounts of soy.
In lay terms, what the MNL formula states is that the CCP depends on three things. The first is the utility. The second is the inherent commonality of characteristics $\phi$: the more lemons there are, the more likely it is that you will buy one. The third is the marginal probability $P$. If you prefer $\xi$ to $\xi^{\prime}$ in 99% of states, then even in the 1% of states where you slightly prefer $\xi^{\prime}$ to $\xi$, you may still end up picking $\xi$ unless you acquire enough information about the state to be quite sure you are in that 1%.
How much each of these three things matters depends on the parameters $\alpha,\lambda$. If $\alpha\lambda$ is large, that means it is harder to learn about the characteristics $\xi$. Thus, $P(\xi|\theta)$ will resemble what would happen if you were to choose randomly, which is represented by the prior $\phi(\xi)$. If $(1-\alpha)\lambda$ is large, that means it is harder to learn about states $\theta$. Thus, $P(\xi|\theta)$ will not differ much from $P(\xi)$, since separating $P(\xi|\theta)$ from $P(\xi)$ means you learned to differentiate $\theta$ from other states. If both are large, then $P(\xi)$ is close to $\phi(\xi)$ and the CCPs are close to both. Conversely, as $\lambda\to 0$, the CCP begins to look like that of a fully-informed, utility (soft-)maximizing agent.
Lastly, section˜5 concerns falsifiable restrictions implied by the results. In particular, entry of products into the market which generate variation in $\phi$ should generate changes in CCP ratios that are not state-dependent. These predictions do not require knowledge of the prior $\phi$ to be tested. However, if the prior is known, the relative cost parameter $\alpha$ is over-identified in the process.
In the finite discrete case, the results may be established by taking Lagrangians. In the remaining sections, we derive the results for the general case, since characteristics may be continuous. The results hereafter are mostly mathematical in nature, whose utility lies in that they provide (hopefully) interesting and (hopefully) novel insights into the problem, as well as connections to other well-known problems. The proofs in these sections are typically deferred to the appendix.
Section˜6 explores the first-step orthogonality property of Shannon entropy-based inattention objectives and its consequences. In brief, the objective (which is a function of the information policy $P$) can be written as an integral in which $P$ appears twice: once in the integrand and once in the form of $dP$. While the product rule suggests that perturbing $P$ ought to affect the objective in two ways, first-step orthogonality establishes that it only locally affects the objective in one: via the re-weighting in the form of $dP$. First-step orthogonality implies that both state-action and state-characteristic have a nice property whereby when the objective is written as a single integral, the integrand is constant.
Section˜7 establishes a connection between rational inattention and entropic optimal transport (EOT). It turns out that the problem of maximizing the state-characteristic objective subject to a second marginal constraint $P|_{\xi}=\nu$ (on top of Bayes plausibility, which is itself a marginal constraint $P|_{\theta}=\mu$) can be re-cast as a Schrödinger bridge problem [7]. The problem comes with a dual which consists of finding Schrödinger potentials $a(\xi),b(\theta)$ which satisfy the associated Schrödinger equations. From EOT theory, we know that the value as a function of the marginal constraint $\nu$ is additively separable and equal to the sum of (the expectations of) the potentials. The potentials have a straightforward Lagrangian interpretation. Loosely speaking, the potential associated with the characteristic $\xi$ corresponds to how much the value would change if we were to change the marginal constraint $\nu$ such that some extra mass were added to $\xi$.
Perhaps the most interesting practical takeaway here is that the optimal CCP subject to two marginal constraints is also a weighted MNL. [4] had established that
$\text{$P$ is optimal}\implies P(\xi|\theta)\propto P(\xi)e^{u(\xi,\theta)/\lambda}$
but this does not necessarily imply that
$\text{$P$ is optimal subject to }P|_{\xi}=\nu\implies P(\xi|\theta)\propto\nu(\xi)e^{u(\xi,\theta)/\lambda}$
Indeed, it is not true. However, the Schrödinger equations imply that $P$ maximizes the objective subject to $P|_{\xi}=\nu,P|_{\theta}=\mu$ if and only if there exists $a,b$ such that
$P(\xi|\theta)=\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}e^{[u(\xi,\theta)-a(\xi)-b(\theta)]/\lambda}$
Thus, $P(\xi|\theta)$ always follows a weighted MNL. Finding such an $a,b$ when $\xi,\theta$ are discrete is a case of the matrix scaling problem. Sinkhorn’s algorithm can be used to find the Schrödinger potentials.
We finish by considering the problem of finding the optimal marginal in section˜8. In [1], the authors write out a concave function over the set of marginals which helps them establish that the Kuhn-Tucker conditions are sufficient, as opposed to only necessary. This function may be obtained by swapping an expectation with a log; by Jensen’s inequality, it upper bounds the objective. It turns out that this function has even nicer properties when $\alpha\in(0,1)$. In particular, it is strictly concave over a convex set and coincides with the objective at their mutual optima. Thus, rather than maximize a more complicated objective, it suffices to solve this much easier convex optimization problem.
Motivation
State-action presents two options for modeling consumer choice with uncertain characteristics: either treat the characteristics as actions – i.e. costlessly observed – or as a state variable, alongside other state variables. The latter option, however, presents an identification problem. It turns out that a DM who does not observe the other state variables can achieve the same payoff at the same information cost as someone who costlessly observes them. Thus, the two behave identically and are observationally equivalent.
Example. Consumers choose between two insurance contracts $j=1,2$. They know one is high-coverage $\xi_{j}=H$ and the other is low-coverage $\xi_{j}=L$ but do not know which is which. The consumer has also two future health expenditure states, $t\in\{H,L\}$, each with probability $1/2$. The consumer receives utility 1 if they choose the ‘correct’ contract (high coverage iff high future expenditures) and 0 otherwise. In other words,
$u(j,\xi_{1},\xi_{2},t)=\textbf{1}\{\xi_{j}=t\}=\begin{cases}1&\xi_{j}=t\\ 0&\xi_{j}\neq t\end{cases}$
Recall the state-action sequence proceeds as follows: the consumer commits to an information policy – a Bayes plausible joint distribution over actions $j$ and states $(\xi_{1},\xi_{2},t)$. Nature then draws an action recommendation and a state. Now, suppose consumer 1 observes $t$ before committing to the information policy
if $\xi_{1}=t$, recommend $j=1$ with probability $p$ and $j=2$ with probability 1p1-p
if $\xi_{2}=t$, recommend $j=2$ with probability $p$ and $j=1$ with probability 1p1-p
and consumer 2 does not observe $t$, but commits to the same information policy. Consider:
  1. 1.
    Do the two consumers have the same information upon receiving the same recommendation $j$?
  2. 2.
    Do the two consumers pay the same information cost?
  3. 3.
    Is the expected payoff (averaged over $t$) the same?
The answer to the first, clearly, is no. Consumer 1 knows $t$ and $j$, whereas consumer 2 only gets $j$. Consumer 1 knows, with probability $p$, that the coverage level will be $t$, whereas consumer 2 merely knows that with probability $p$, the coverage level will be appropriate for their future health expenditures – but does not know what either are. The answer to the second is yes. The answer to the third is also yes – for both, their utility will be 1 with probability $p$ and 0 with 1p1-p.
Ergo, consumer 2 is able to achieve the same expected payoff at the same information cost, despite having less information before committing to the information policy and after receiving the recommendation. Worse, for the econometrician, the conditional choice probability for both consumers is the same, so they are indistinguishable.
What is happening is that both consumers are acquiring the same amount of information about whether $j$ is the “right” choice. Consumer 1 already knows $t$ and is doing so by acquiring information about what $j$ is. But consumer 2 is doing so without acquiring any information about what $j$ is or why it might be the right choice. What the state-action model accounts for is the fact that for each $t$, the probability of choosing $\xi_{j}=t$ differs from $1/2$ – what it would be if no information was acquired; what it doesn’t account for – or rather, treats as costless – is the information about $t$ itself. This issue can be formalized as follows.
Observational equivalence in state-action models when characteristics are state variables. In general, could an econometrician determine whether an insurance buyer in the state-action model has a priori private information about their future health expenditures $\theta$? Consider the canonical consumer problem setting (see e.g. [1]), the action space is $A=\{1,...,J\}$, and the utility vector is $\textbf{u}(\boldsymbol{\xi},\theta)=(u(1,\boldsymbol{\xi},\theta),...,u(J,\boldsymbol{\xi},\theta))$, which is a function of the random variables $j\in A$, $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$, and $\theta\in\Theta$. The state-action model predicts that the optimal conditional choice probability is
$P(j|\boldsymbol{\xi},\theta)=P(j|\textbf{u}(\boldsymbol{\xi},\theta))=\frac{e^{u(j,\boldsymbol{\xi},\theta)}}{\sum_{k\in[J]}e^{u(k,\boldsymbol{\xi},\theta)}}$
This equation holds regardless of whether $\theta$ is assumed to be known to the consumer a priori or not. Thus, whether $\theta$ is known a priori is not identified in the state-action model.
Mathematically, we observe that by the law of iterated expectations,
$I_{P}(j,\textbf{u}):=\int\log\left(\frac{dP(j,\textbf{u})}{dP(j)dP(\textbf{u)}}\right)\>dP=\iint\log\left(\frac{dP(j,\textbf{u}|\theta)}{dP(j|\theta)dP(\textbf{u}|\theta)}\right)\>dP(j,\textbf{u}|\theta)\>dP(\theta)=\int I_{P}(j|_{\theta},\textbf{u}|_{\theta})\>dP(\theta)$
because $P(j|\theta)=P(j)$. So, the DM maximizes
$E_{P}[u]-\lambda I_{P}(j,\textbf{u})\equiv E_{P}\bigg{[}E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})\bigg{]}$
if and only if they maximize $E_{P}[u|\theta]-\lambda I_{P}(j|_{\theta},\textbf{u}|_{\theta})$ for almost every $\theta$. In words, the DM solves the state-action rational inattention problem overall if and only if they solve the state-action rational inattention problem in every health state $\theta$.
One final observation is that $I_{P}(j|_{\theta},\textbf{u}|_{\theta})=D_{KL}(P(\xi|\theta)\|\phi(\xi))$ for a natural prior $\phi$. To see this, suppose that out of the $J$ options, $k_{x}$ of them had the characteristic $x\in X$. A natural prior, then, would be $\phi(x)=k_{x}/J$. Without loss, suppose $\xi_{j}=x$. Then, $P(\xi_{j}|\theta,\boldsymbol{\xi})=P(x|\theta)/k_{x}$. Thus,
$\frac{P(j|\textbf{u})}{P(j)}=\frac{P(j|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(\xi_{j}|\theta,\boldsymbol{\xi})}{1/J}=\frac{P(x|\theta)/k_{x}}{1/J}=\frac{P(x|\theta)}{\phi(x)}$
The need for consumer choice uncertainty. Given the non-identifiability problem, it may be tempting to simply assume that a consumer could observe the characteristics of their choices costlessly, and simply use the state-action model to model the uncertainty generated by $\theta$. Indeed, this is what most empirical discrete choice models implicitly do when $\xi$ is taken to be the choice variable. This is reasonable since the econometrician observes $P(\xi,\theta)$, not $P(k,\textbf{u})$. However, there are many circumstances in which the state-characteristic model we propose herein may be helpful.
For one, the choice probability predicted by the state-action model may be non-unique. Second, we may want to incorporate information that some options are more likely than others. Third, we may observe choices which appear strictly dominated, as is empirically common [2], which cannot be explained by modeling characteristics as actions in the state-action model. Fourth, there is no incorporation of the inherent ubiquity of certain options.
To illustrate these problems in structural econometric context, suppose Harold Zurcher is buying used buses, and $X=\{\text{red bus, blue bus, lemon bus}\}$ such that $\text{red}=\text{blue}\succ\text{lemon}$ in all states. In the state-action model without choice uncertainty, he would ever choose a lemon, no matter how abundant they were. And even if we knew red buses outnumbered blue buses 100-to-1, the state-action model does not prescribe a unique optimum policy – it could be optimal for Zurcher to only buy blue (or red).
So what are we missing? What we have shown is that when characteristics are modeled as state variables, the state-action imposes a cost function which is effectively $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$, and when characteristics are modeled as actions, then naturally the cost function is $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$. In the former, the DM behaves as though observing $\theta$ is free. In the latter, the DM observes $\xi$ for free.
To wit, the natural extension of the state-action model which we propose in the next section contains two terms: $E_{P}[D_{KL}(P(\xi|\theta)\|\phi(\xi))]$ and $I_{P}(\xi,\theta)=E_{P}[D_{KL}(P(\theta|\xi)\|\mu(\theta))]$.
Model
Notation. Let $m,m^{\prime}$ be measures defined on the same measurable space. $m\ll m^{\prime}$ denotes absolute continuity. $m\ll m^{\prime}\ll m$ is interchangeable with $\sim m^{\prime}. m=m$$m^{\prime}$ if $m(A)=m^{\prime}(A)$ for each measurable set. The KL-divergence between two probabilities is defined
$D_{KL}(m\|m^{\prime})=\begin{cases}\int\log\left(\frac{dm}{dm^{\prime}}\right)\>dm&(m\ll m^{\prime})\\ \infty&\text{else}\end{cases}$
By convention, $\infty\cdot 0=0$ for integration purposes.222see [3]
Model. Fix an ambient probability space $(\Omega,{\cal F},\mathbb{P})$, a space of characteristics $X$ with an exogenous prior probability $\phi\in\Delta X$, and a space of states $\Theta$ with prior probability $\mu\in\Delta\Theta$. The decision-maker’s (DM) problem is to maximize expected utility $u$ less an information cost $\kappa_{\alpha,\lambda}$
$\sup_{P\in\Delta(X\times\Theta)}U(P):=E_{P}[u(\xi,\theta)]-\kappa_{\alpha,\lambda}(P)$ (1)
subject to
  1. 1.
    (Bayes plausibility) $P|_{\theta}=\mu$
  2. 2.
    (marginal feasibility) $P|_{\xi}\ll\phi$
The cost function $\kappa_{\alpha,\lambda}$ can be written in several ways:
[…]
$\equiv\alpha\lambda\bigg{(}\underbrace{D_{KL}(P(\xi)\|\phi(\xi))+I_{P}(\xi,\theta)}_{\text{information about }\xi}\bigg{)}+(1-\alpha)\lambda\underbrace{I_{P}(\xi,\theta)}_{\text{information about $\theta$}}$
$\equiv\alpha\lambda\underbrace{D_{KL}(P(\xi)\|\phi(\xi))}_{\text{``vertical information''}}+\lambda\underbrace{I_{P}(\xi,\theta)}_{\text{``horizontal information''}}$
where $I_{P}$ denotes mutual information under $P$. To generalize, we impose further regularity conditions:
  1. 1.
    $X,\Theta$ are Polish spaces with Borel $\sigma$-fields ${\cal B}_{X},{\cal B}_{\Theta}$, over which $\phi$ and $\mu$ are respectively defined
  2. 2.
    $u:X\times\Theta\to[\underline{u},\overline{u}]$ is real-valued, bounded, and measurable
Example: state-action with costlessly observed $\xi$. When $\alpha=0$, $\kappa_{\alpha,\lambda}(P)$ simplifies to $I_{P}(\xi,\theta)$, which corresponds to the state-action rational inattention model. Id est, when $\alpha=0$, the DM can choose characteristics directly, just as in the state-action model, the DM chooses actions directly.
Example: state-action with costlessly observed $\theta$. Consider a state-action model in which the DM knows their type $\theta$ a priori. The DM has an action set $A=\{1,...,J\}$ consisting of $J$ products with $J$ characteristics. The DM knows what the characteristics are, but does not know how they are permuted among the $J$ actions. Let $\boldsymbol{\xi}=(\xi_{1},...,\xi_{J})$ denote these characteristics. Denote $\xi$ with the characteristic of the product he eventually chooses. The mutual information $I_{P}(j,\boldsymbol{\xi})$ between action $j\in[J]$ and $\boldsymbol{\xi}$ is
$\int D_{KL}(P(\boldsymbol{\xi}|j)\|P(\boldsymbol{\xi}))\>dP(j)=D_{KL}(P(\boldsymbol{\xi}|j=1)\|P(\boldsymbol{\xi}))=D_{KL}(P(\xi_{1}|j=1)\|P(\xi_{1}))=D_{KL}(P(\xi)\|\phi(\xi))$
The third equality follows from the observation that choosing $j=1$ provides only information about what $\xi_{1}$ and is not (meaning it provides no information about $(\xi_{2},...,\xi_{J})$ beyond the fact that they are not $\xi_{1}$). Thus, this corresponds to the state-characteristic model with $\alpha=1$ and $\phi(\xi)=1/J$.
Example: Maxwell-Boltzmann statistics. When $\alpha=1$, the law of iterated expectations yields
$U(P)$ $=E_{\theta\sim\mu}\bigg{[}E_{P}\bigg{[}u(\xi,\theta)-\lambda D_{KL}(P(\xi|\theta)\|\phi(\xi))\ |\ \theta\bigg{]}\bigg{]}$
$=E_{\theta\sim\mu}\bigg{[}E_{P}[u(\xi,\theta)|\theta]-\lambda\cdot E_{\phi}\bigg{[}\frac{dP(\xi|\theta)}{d\phi(\xi)}\log\left(\frac{dP(\xi|\theta)}{d\phi(\xi)}\right)\bigg{]}\bigg{]}$
That is, $U(P)$ is maximized iff the conditional distribution maximizes what is inside the expectation for every $\theta$. Independent of the forthcoming results, the Donsker-Varadhan variational formula implies the conditional choice probability follows Maxwell-Boltzmann statistics:
$P(\xi|\theta)=\frac{\phi(\xi)e^{u(\xi,\theta)/\lambda}}{Z(\theta;P)}$
From a physical view, the utility maximization problem can be formulated as a free energy minimization problem. The prior $\phi$ corresponds to the “degeneracy” of the energy level. The utility corresponds to the enthalpy. $\lambda$ corresponds to the temperature. $E_{\phi}\bigg{[}\frac{dP(\xi|\theta)}{d\phi(\xi)}\log\left(\frac{dP(\xi|\theta)}{d\phi(\xi)}\right)\bigg{]}$ is the conditional differential entropy (the cross-entropy between $P(\cdot|\theta)$ and $\phi$).
Measure-theoretic notes. Some care must be taken when considering the distinction between a function $f$ and its associated equivalence class of functions $g$ for which $f=g$ a.e. In general, the ‘choice’ of $P$ does not precisely pin down its disintegration kernels or associated densities. Though not fatal, the ambiguity of possibly infinite densities clutters analysis with qualifiers. To simplify things, we restrict the DM to choosing disintegrations and densities which satisfy
$\frac{dP(\xi,\theta)}{d(P|_{\xi}\otimes\mu)(\xi,\theta)}=\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}<\infty$ $\frac{dP(\xi)}{d\phi(\xi)}<\infty$ $\frac{dP(\theta)}{d\mu(\theta)}=1$ (2)
everywhere, not just almost everywhere. The existence of these are established in the following lemmas.
Lemma 1.
If $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$, then $P(\cdot|\theta)\ll\phi$ $\mu$-a.s. and $P\ll\phi\otimes\mu$.
Lemma 2.
Let $P|_{\xi}\ll\phi$ and $P|_{\theta}=\mu$. There exist valid densities such that eq.˜2 is satisfied.
The proofs are deferred to the appendix.
Main Results
We consider the case where $\alpha\in(0,1)$. We normalize $\lambda$ to $1$, since maximizing $U(P)$ is the same as maximizing $U(P)/\lambda$. To obtain results for $\lambda\neq 1$, simply take all results herein in which $u(\xi,\theta)$ and associated additive terms appear and divide them by $\lambda$.
Theorem 3 (Gibbs property).
If $P$ solves the objective eq.˜1 then
$\xi\mapsto Y(\xi,\theta;P)$ is constant $P(\cdot|\theta)$-a.s. for $\mu$-a.e. $\theta$
Corollary 3.1 (Weighted multinomial logit).
If $P$ solves the objective eq.˜1 then
$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(\xi,\theta)}}{Z(\theta;P)}$ (3)
for some normalizing partition function $Z$.
Theorem 4.
The optimal marginal $P|_{\xi}$ is unique, and therefore the optimal CCP is unique.
Theorem 5.
If $P$ maximizes $V$, then $P|_{\xi}\ll\phi\ll P|_{\xi}$.
These follow from corollaries˜7.1, 8 and 15.1. In the discrete case, these can be derived by taking Lagrangians.
Testable restrictions
In the context of discrete-choice, the model provides falsifiable restrictions which describe how the CCPs should react to the entry of a new product. It is not necessary that the econometrician knows $\phi$, but knowing $\phi$ means $\alpha$ is (over-)identified.
Abstractly, $\phi$ is defined as the choice probability when no information is acquired. A natural assumption is to take $\phi$ to be 1 divided by the number of characteristics available in the market, but $\phi$ could also be other things. In markets where people who acquire little to no information may be identified, $\phi$ is identified.
(Over-)Identification of $\alpha$ via entry. Assume the market has $J$ characteristics $\xi_{1},...,\xi_{J}$ and $\phi$ is known. Consider entry by a new product which changes $\phi$ to $\phi^{\prime}$, which induces a change in the marginal $P$ to $P^{\prime}$. For any $\theta$, we have that
$\frac{P(\xi|\theta)}{P^{\prime}(\xi|\theta)}=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}}{\phi^{\prime}(\xi)^{\alpha}P^{\prime}(\xi)^{1-\alpha}}\frac{Z^{\prime}(\theta;P^{\prime}|_{\xi})}{Z(\theta;P|_{\xi})}$
For any pair $\xi_{1},\xi_{2}\in X$, we have
$\frac{P(\xi_{1}|\theta)}{P^{\prime}(\xi_{1}|\theta)}\frac{P^{\prime}(\xi_{2}|\theta)}{P(\xi_{2}|\theta)}=\frac{\phi(\xi_{1})^{\alpha}P(\xi_{1})^{1-\alpha}}{\phi^{\prime}(\xi_{1})^{\alpha}P^{\prime}(\xi_{1})^{1-\alpha}}\frac{\phi^{\prime}(\xi_{2})^{\alpha}P^{\prime}(\xi_{2})^{1-\alpha}}{\phi(\xi_{2})^{\alpha}P(\xi_{2})^{1-\alpha}}$
That is, an over-identifying feature of the model is that we can construct a ratio that depends on $\theta$ on the left-hand side but not on the right-hand side. We can test whether it is true that for all pairs of $\xi$, the left-hand side is constant w.r.t. $\theta$. If this is not true, the model is easily rejected. What’s more, testing this restriction does not require any knowledge of $\phi$, which does not appear on the left-hand side. However, knowing $\phi$ naturally pins down $\alpha$.
First-Step Orthogonality and Its Implications
First-step orthogonality is a property of moments of the form $\int f_{P}(\omega)\>dP(\omega)$, where the integrand $f$ depends on the probability with respect to which we are integrating. Loosely, it states that a perturbation in $P$ locally affects the moment only through the re-weighting $dP$. It constitutes a sufficient condition for the Gibbs property and additive separability. It is a relatively unique property of entropy-based costs which essentially comes from the fact that $\frac{d}{dx}\log x=1/x$.
Notation. Let $\cal P$ be the set of Bayes-plausible feasible probabilities. Let $H\in{\cal P}$. Define a path $P_{\varepsilon}:=(1-\varepsilon)P+\varepsilon H$. Clearly, $P_{\varepsilon}$ is defined on $\varepsilon\in[l,r]$ for some $l\leq 0$ and $r\geq 1$, and $P_{\varepsilon}\subset{\cal P}$.
Lemma 6 (First-step orthogonality).
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}E_{P_{t}}[Y(\xi,\theta;P_{\varepsilon})]=0$ for $t\in[l,r)$
Theorem 7 (First-step orthogonality II).
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}U(P_{\varepsilon})=E_{H}[Y(\xi,\theta;P_{t})]-E_{P}[Y(\xi,\theta;P_{t})]$ for $t\in[l,r)$
Corollary 7.1 (Gibbs property).
If $P$ maximizes $U$ then $\xi\mapsto Y(\xi,\theta;P)$ is constant a.s.
Proof.
We have $\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}U(P_{\varepsilon})=E_{\theta\sim\mu}\left[\int Y(\xi,\theta;P)\>d(H-P)(\xi|\theta)\right]$. For every $\theta$ on which the property is not satisfied, we can take $H(\xi|\theta)$ to place mass only on the points where $Y$ is largest such that $\int Y(\xi,\theta;P)\>d(H-P)(\xi|\theta)>0$. ∎
By the same token if $P$ maximizes $U$ subject to a marginal constraint $P|_{\xi}=\nu$, then $Y$ should be additively separable. More on this later.
Lemma 8.
Let $\alpha>0$. If $P$ maximizes $U$ then the density $\frac{dP(\xi)}{d\phi(\xi)}$ does not get arbitrarily close to 0 and does not become arbitrarily large.
Schrödinger Potentials
The problem faced by the DM can be separated into two problems: i) maximizing the objective eq.˜1 subject to a further marginal constraint $P|_{\xi}=\nu$ and ii) solving for the optimal marginal $\nu$. We observe here that the former is an instance of the entropic optimal transport problem. The constrained optimal marginal is called the Schrödinger bridge from $\nu$ to $\mu$.
In lemma˜8, we established that if $P$ maximizes $U$ subject to feasibility and Bayes plausibility, then $P|_{\xi}\sim\phi$ and $dP(\xi)/d\phi(\xi)$ is upper bounded and lower bounded away from zero. So in this section, we assume $P|_{\xi}$ is restricted to a marginal $\nu$ for which $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from zero.
Define ${\cal Q}(\nu)=\{P\in\Delta(X\times\Theta):P|_{\xi}=\nu,P|_{\theta}=\mu\}$. We let
$V(\nu)=\sup_{P\in{\cal Q}(\nu)}U(P)$
denote the value of imposing the constraint $P|_{\xi}=\nu$.
We can rewrite U(P)-U(P) as:
$-U(P)$ $=\int-{u(\xi,\theta)}+\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)\>dP+\int\log\left(\frac{dP}{d(\nu\otimes\mu)}\right)\>dP$
and turn to minimizing
$-V(\nu)=\inf_{P\in{\cal Q}(\nu)}\int c(\xi,\theta)\>dP-D_{KL}(P\|\nu\otimes\mu)$ (4)
where $c(\xi,\theta)=-{u(\xi,\theta)}+\alpha\log\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)\in L^{\infty}(\nu\otimes\mu)$.
This is called the Entropic Optimal Transport (EOT) problem. [^3] For simplicity, we may assume $c(\xi,\theta)$ is totally bounded, since the worse case scenario choice of $\nu$ can be set to $\phi$, in which case c=uc=-u, which we assumed to be totally bounded. The following theorems concern properties of EOT and can be found in [6], up to notational changes.
Theorem 9 (Existence and uniqueness of potentials).
If $P$ solves eq.˜4, then there exist functions $a_{\nu}\in L^{1}(\nu),-b_{\nu}\in L^{1}(\mu)$ such that
$\frac{dP}{d(\nu\otimes\mu)}=e^{-a_{\nu}(\xi)-b_{\nu}(\theta)-c(\xi,\theta)}=e^{u(\xi,\theta)-a_{\nu}(\xi)-b_{\nu}(\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}$ (5)
These are called the Schrödinger potentials. Conversely, if there exists potentials such that eq.˜5 holds, then $P$ must be optimal. Moreover, potentials are unique up to translation. [^4] That is, for any two pairs of potentials $(a_{\nu},b_{\nu})$ and $(a_{\nu}^{\prime},b_{\nu}^{\prime})$, we have $a_{\nu}-a^{\prime}_{\nu}=b_{\nu}^{\prime}-b_{\nu}$.
Corollary 9.1.
If $P$ solves eq.˜4 then $Y(\xi,\theta;P)=a_{\nu}(\xi)+b_{\nu}(\theta)$ and so $U(P)=E_{\nu}[a_{\nu}(\xi)]+E_{\mu}[b_{\nu}(\theta)]$
Theorem 10 (Schrödinger Equations).
$a_{\nu}$ and $b_{\nu}$ constitute a pair of Schrödinger potentials if and only if they satisfy
$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{e^{b_{\nu}(\theta)}}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}}\>d\nu(\xi)$
Ergo,
$e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{a_{\nu}(\xi)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{\alpha}}\>d\nu(\xi)}\>d\mu(\theta)$ $e^{b_{\nu}(\theta)}=\int\frac{e^{u(\xi,\theta)}}{\int\frac{e^{u(\xi,\theta)}}{e^{b_{\nu}(\theta)}}\>d\mu(\theta)}\>d\nu(\xi)$
Theorem 11 (Duality).
$V(\nu):=\sup_{P:\ P|_{\xi}=\nu,P|_{\theta}=\mu}U(P)$
$V(\nu)$ $=\inf_{a\in L^{1}(\nu),\ b\in L^{1}(\mu)}\int a\>d\nu+\int b\>d\mu+\iint e^{-a-b-c}\>d\nu\>d\mu-1$
$=\inf_{a\in L^{1}(\nu),\ b\in L^{1}(\mu)}\int a\>d\nu+\int b\>d\mu+\iint e^{u(\xi,\theta)-a-b}\>d\phi^{\alpha}d\nu^{1-\alpha}\>d\mu-1$
Note that because $d\nu/d\phi$ and $d\phi/d\nu$ are bounded away from 0, by Hölder’s equality with coefficients $p=1$ and $q=\infty$, we can interchange $a\in L^{1}(\nu)$ with $a\in L^{1}(\phi)$.
Corollary 11.1.
Consider a path $\nu_{\varepsilon}$ such that at $\varepsilon=0$, $\varepsilon\mapsto\frac{d\nu_{\varepsilon}(\xi)}{d\phi(\xi)}$ is continuous for $\phi$-a.e. $\xi$, and $d\nu_{\varepsilon}/d\phi$ is uniformly bounded away from 0 and $\infty$ on $\varepsilon\in(-r,r)$. Then $\varepsilon\mapsto V(\nu_{\varepsilon})$ is continuous at 0.
In the discrete case, the Schrödinger potentials can be viewed as the Lagrange multipliers for the marginal constraints. In other words, $a_{\nu}(x)$ can be viewed as the rate of increase in $V$ when adding an infinitessimal mass of $x$ to $\nu$. This is because
$V(\nu)=\int a_{\nu}\>d\nu+\int b_{\nu}\>d\mu$
so in the discrete case, applying the Envelope theorem and normalizing $\int a_{\nu}\>d\nu=0$ gives
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}V\bigg{(}(1-\varepsilon)\nu+\varepsilon\delta_{x}\bigg{)}=a_{\nu}(x)$
The continuous case is more involved and discussion is deferred to the appendix.
The Jensen Upper Bound of the Marginal Value
If one writes out the objective – a function of the joint $P$ – swaps the log with the first integral, one obtains a function of the marginal $P|_{\xi}$. When $\alpha\in(0,1)$ this function is strictly convex, and, by Jensen’s Inequality, upper bounds $V$. We show that their maxima coincide, meaning that solving for the optimal marginal is a relatively simple convex optimization problem.
Define $f(\nu)=E_{\theta\sim\mu}\left[\log\left(E_{\xi\sim\phi}\left[e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}\right]\right)\right]\equiv E_{\theta\sim\mu}[\log(Z(\theta;\nu)]$
Theorem 12.
$f(P|_{\xi})\geq U(P)$
Theorem 13.
$f$ is strictly concave when $\alpha\in(0,1)$ and therefore has a unique maximum.
Theorem 14.
If $P$ maximizes $U$ then $f(P|_{\xi})=U(P)$.
Proof.
This follows from plugging in the weighted MNL formula. ∎
Theorem 15.
The maximum of $f$ and $U$ coincide. That is, if $\nu$ maximizes $f$, then there exists a $P$ with $P|_{\xi}=\nu$ which maximizes $U$.
Corollary 15.1.
The maximum of $U$ is unique.
Conclusion
We study a model of rational inattention to states $\theta$ and hedonic characteristics $\xi$. The DM pays the standard linear-in-Shannon cost for the mutual information contained in the joint distribution over $(\xi,\theta)$. In addition, some choices are intrinsically more common, as represented by the prior $\phi$, and so the DM pays a cost for diverging from $\phi$.
The DM’s conditional choice probability satisfies a weighted multinomial logit:
$P(\xi|\theta)=\frac{\phi(\xi)^{\alpha}P(\xi)^{1-\alpha}e^{u(\xi,\theta)}}{Z(\theta;P)}$
In the state-action model, the DM has full information about their actions and is only learning about the state, and so narrows down the set of choices to a “consideration set.” Because there are many possible optimal choices for marginals $P(a)$, there may not be a unique solution. By contrast, in our model, the DM does not observe the hedonic characteristics of their selections; because it gets increasingly costly to rule out characteristics with increasing certainty, the DM never rules out any choice. Thus, the marginal lies in the interior and, because of strict concavity, is unique.
Appendix
Omitted proofs
Proof of lemma˜1.
Let $P(\cdot|\theta)$ be any disintegration kernel. To see that $P(\cdot|\theta)\ll\phi$, $\mu$-a.s., let $\phi(A)=0$. Then,
$P|_{\xi}(A)=\int P(A|\theta)\>d\mu(\theta)=0$
which implies that $P(A|\theta)=0$ $\theta$-a.s. Therefore, $dP(\xi|\theta)/d\phi(\xi)<\infty$ $\phi\otimes\mu$-a.e. To see $P\ll\phi\otimes\mu$, let $(\phi\otimes\mu)(S)=0$. Then,
$P(S)=\int\textbf{1}_{S}\>dP=\iint\textbf{1}_{S}\frac{dP(\xi|\theta)}{d\phi(\xi)}\>d\phi(\xi)\>d\mu(\theta)$
$\textbf{1}_{S}=0$ $\phi\otimes\mu$-a.e., so $P(S)=0$. So, $dP/d(\phi\otimes\mu)<\infty$, $\phi\otimes\mu$-a.e. ∎
Proof of lemma˜2.
Since $P\ll\phi\otimes\mu$, by the Radon-Nikodym theorem, there is an equivalence class of a.e.-equal densities which are a.e.-finite. By construction, no matter which element of the equivalence class one takes,
$\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\mu(\theta)<\infty\text{; $\phi$-a.s.}$ $\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\phi(\xi)=1\text{; $\mu$-a.s.}$ $\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}<\infty\text{; $\phi\otimes\mu$-a.s.}$
so on the null set of points and fibers where these almost-sure properties do not hold, we can re-define $\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}$ to 1, so that the properties now hold everywhere, not just a.e. Since we are changing the function on a null set, it is still in the equivalence class. Then, we define the other densities as follows:
$\frac{dP(\xi|\theta)}{d\phi(\xi)}=\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}$ $\frac{dP(\xi)}{d\phi(\xi)}=\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\mu(\theta)$ $\frac{dP(\theta)}{d\mu(\theta)}=\int\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}\>d\phi(\xi)$
and
$\frac{dP(\xi|\theta)}{dP(\xi)}=\frac{dP(\theta|\xi)}{d\mu(\theta)}=\begin{cases}\frac{\frac{dP(\xi,\theta)}{d(\phi\otimes\mu)(\xi,\theta)}}{\int\frac{dP(\xi,t)}{d(\phi\otimes\mu)(\xi,t)}\>d\mu(t)}&\text{if the denominator is positive}\\ 1&\text{otherwise}\end{cases}$
It can then be seen that the kernels defined $P(A|\theta):=\int_{A}\frac{dP(\xi|\theta)}{d\phi(\xi)}\>d\phi(\xi)$ and $P(B|\xi)=\int_{B}\frac{dP(\theta|\xi)}{d\mu(\theta)}\>d\mu(\theta)$ are valid disintegration kernels. ∎
Proof of lemma˜6.
We have
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}\frac{dP_{t}(\theta|\xi)}{d\mu(\theta)}\log\left(\frac{dP_{\varepsilon}(\theta|\xi)}{d\mu(\theta)}\right)=\frac{d(H-P)(\theta|\xi)}{d\mu(\theta)}\in L^{1}(\mu)$
for all $t\in[0,1]$. The integral of which is
$\iint\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}\frac{dP_{t}(\theta|\xi)}{d\mu(\theta)}\log\left(\frac{dP_{\varepsilon}(\theta|\xi)}{d\mu(\theta)}\right)\>d\mu(\theta)\>dP_{t}(\xi)=\iint\frac{d(H-P)(\theta|\xi)}{d\mu(\theta)}\>d\mu(\theta)\>dP_{t}(\xi)=0$
From a similar calculation, one obtains
$\int\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}\frac{dP_{\varepsilon}(\xi)}{d\phi(\xi)}\log\left(\frac{dP_{\varepsilon}(\xi)}{d\phi(\xi)}\right)\>d\phi(\xi)=0$
The exchangeability of the integral and the derivative follows from [3, Theorem 2.27]; in particular, the sufficient condition is that the absolute value of the derivative of the integrand is bounded by an element of $L^{1}$. The result follows by construction of $Y$. ∎
Proof of theorem˜7.
By the product rule,
$\displaystyle\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}U(P_{\varepsilon})$ $=\lim_{\varepsilon\downarrow t}\frac{1}{\varepsilon-t}\left[\int Y(\xi,\theta;P_{\varepsilon})\>dP_{\varepsilon}-\int Y(\xi,\theta;P_{t})\>dP_{t}\right]$
$=\lim_{\varepsilon\downarrow t}\frac{1}{\varepsilon-t}\left[\int Y(\xi,\theta;P_{\varepsilon})\>dP_{\varepsilon}-\int Y(\xi,\theta;P_{\varepsilon})\>dP_{t}+\int Y(\xi,\theta;P_{\varepsilon})\>dP_{t}-\int Y(\xi,\theta;P_{t})\>dP_{t}\right]$
$=\left(\lim_{\varepsilon\downarrow t}\frac{1}{\varepsilon-t}\left[\int Y(\xi,\theta;P_{\varepsilon})\>dP_{\varepsilon}-\int Y(\xi,\theta;P_{\varepsilon})\>dP_{t}\right]\right)+\underbrace{\displaystyle\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t^{+}}E_{P_{t}}\left[Y(\xi,\theta;P_{\varepsilon})\right]}_{=0}$
$=\lim_{\varepsilon\downarrow t}\int Y(\xi,\theta;P_{\varepsilon})\>d(H-P)=\int Y(\xi,\theta;P_{t})\>d(H-P)$
Proof of lemma˜8.
Let $\underline{u}\leq u(\xi,\theta)\leq\overline{u}$. For contradiction, suppose for every $M\in\mathbb{R}$, there was some set $A_{M}$ of mass $\phi(A_{M})>0$ where $\alpha\log\left(dP(\xi)/d\phi(\xi)\right)>M$. Then, since $Y(\xi,\theta;P)=b(\theta)$, $P$-a.s., then it must be the case that for $P(\cdot|\theta)$-a.e. $\xi\in A_{M}$,
$b(\theta)=u(\xi,\theta)-\alpha\log\left(\frac{dP(\xi)}{d\phi(\xi)}\right)-\log\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)>M+u(\xi,\theta)-\log\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)$
meaning for $P$-a.e. $(\xi,\theta)$,
$\frac{dP(\xi|\theta)}{dP(\xi)}>e^{M+u(\xi,\theta)-b(\theta)}$
which implies
$1=\int\frac{dP(\xi|\theta)}{dP(\xi)}\>d\mu(\theta)>e^{M}\int e^{u(\xi,\theta)-b(\theta)}\>d\mu(\theta)$
for every $M$. This implies that $\int e^{-b(\theta)}\;d\mu(\theta)=0$ thus raising the contradiction.
Along the same lines, suppose for every $M\in\mathbb{R}$, there was some set on which $\alpha\log\left(dP(\xi)/d\phi(\xi)\right)<-M$. We have $b(\theta)<-M+u(\xi,\theta)-\log\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)$, and thus
$1=\int\frac{dP(\xi|\theta)}{dP(\xi)}\>dP(\xi)<e^{-M}e^{b(\theta)}\int e^{u(\xi,\theta)}\>dP(\xi)\leq e^{-M}e^{b(\theta)}e^{\overline{u}}$
so $e^{b(\theta)}=\infty$ everywhere, which also raises a contradiction. ∎
Proof of corollary˜11.1.
Suppose $\gamma,\nu$ are two marginals, and without loss $V(\gamma)\geq V(\nu)$. We can upper bound $V(\gamma)$ by using the “wrong” Schrödinger potentials in the dual objective. In particular, let $\gamma=\nu_{\varepsilon}$
$|V(\nu_{\varepsilon})-V(\nu)|\leq\bigg{|}\int a_{\nu}\>d(\nu_{\varepsilon}-\nu)+\iint e^{u(\xi,\theta)-a_{\nu}-b_{\nu}}\left[\left(\frac{d\nu_{\varepsilon}}{d\phi}\right)^{1-\alpha}-\left(\frac{d\nu}{d\phi}\right)^{1-\alpha}\right]\>d\phi\>d\mu-1\bigg{|}$
But this is continuous w.r.t. $\varepsilon$. Since $\frac{d\nu_{\varepsilon}}{d\phi}$ is uniformly bounded away from 0 and $\infty$ in a neighbourhood of $\varepsilon=0$, we apply the Dominated Convergence Theorem to get $V(\nu_{\varepsilon})\to V(\nu)$. ∎
Proof of theorem˜12.
By Jensen’s inequality,
$U(P)$ $=E_{\theta\sim\mu}\left[\int\log\left(e^{u(\xi,\theta)}\left(\frac{dP(\xi)}{d\phi(\xi)}\right)^{-\alpha}\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)^{-1}\right)dP(\xi|\theta)\right]$
$\leq E_{\theta\sim\mu}\left[\log\left(\int e^{u(\xi,\theta)}\left(\frac{dP(\xi)}{d\phi(\xi)}\right)^{-\alpha}\left(\frac{dP(\xi|\theta)}{dP(\xi)}\right)^{-1}dP(\xi|\theta)\right)\right]$ (Jensen’s Inequality)
$=E_{\theta\sim\mu}\left[\log\left(\int e^{u(\xi,\theta)}\left(\frac{dP(\xi)}{d\phi(\xi)}\right)^{1-\alpha}d\phi(\xi)\right)\right]$ (change of measure)
$=f(P|_{\xi})$
Proof of theorem˜13.
Let $\gamma,\nu$ be feasible probability measures and $\beta\in(0,1)$. Since $\alpha\in(0,1)$, $x\mapsto x^{1-\alpha}$ is strictly concave, it follows
$Z(\theta;\beta\gamma+(1-\beta)\nu)$ $=\int\left(\beta\frac{d\gamma(\xi)}{d\phi(\xi)}+(1-\beta)\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}e^{u(\xi,\theta)/\lambda}d\phi(\xi)$
$>\beta\int\left(\frac{d\gamma(\xi)}{d\phi(\xi)}\right)^{1-\alpha}e^{u(\xi,\theta)/\lambda}d\phi(\xi)+(1-\beta)\int\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{1-\alpha}e^{u(\xi,\theta)/\lambda}d\phi(\xi)$
$=\beta Z(\theta;\gamma)+(1-\beta)Z(\theta;\nu)$
Moreover, $\log$ is strictly concave and strictly increasing, so
$\log\left(Z(\theta;\beta\gamma+(1-\beta)\nu)\right)>\log\left(\beta Z(\theta;\gamma)+(1-\beta)Z(\theta;\nu)\right)>\beta\log(Z(\theta;\gamma))+(1-\beta)\log(Z(\theta,\nu))$
From which we conclude that
$f\left(\beta\gamma+(1-\beta)\nu\right)$ $=\int\log\left(Z(\theta;\beta\gamma+(1-\beta)\nu)\right)\>d\mu(\theta)$
$>\beta\int\log(Z(\theta;\gamma))\>d\mu(\theta)+(1-\beta)\int\log(Z(\theta;\nu))\>d\mu(\theta)$
$=\beta f\left(\gamma\right)+(1-\beta)f\left(\nu\right)$
as needed. ∎
Proof of theorem˜14.
If $\nu$ maximizes $f$, then a first-order condition is that for any $\gamma$ which defines a path $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$, we have $df(\nu_{\varepsilon})/d\varepsilon\leq 0$ at 0, which implies:
$0$ $\geq\frac{df(\nu_{\varepsilon})}{d\varepsilon}\bigg{|}_{\varepsilon=0}\geq\int\frac{\int(1-\alpha)e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\frac{d(\gamma-\nu)(\xi)}{d\phi(\xi)}d\phi(\xi)}{\int e^{u(\xi,\theta)}d\phi^{\alpha}d\nu^{1-\alpha}}d\mu=\int\frac{\int e^{u(x,\theta)}\left(\frac{d\nu(x)}{d\phi(x)}\right)^{-\alpha}\>d\gamma(x)}{\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\>d\nu(\xi)}-1\>d\mu$
Since this holds for all $\gamma\ll\phi$, it follows for $\phi$-a.e. $x$,
$\int\frac{e^{u(x,\theta)}\left(\frac{d\nu(x)}{d\phi(x)}\right)^{-\alpha}}{\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\>d\nu(\xi)}-1\>d\mu\leq 0$
(To show this, assume for contradiction that there is a set $A$ with $\phi(A)>0$ such that the converse is true. Then, take $\gamma$ such that $d\gamma(x)/d\phi(x)>0$ iff $x\in A$. Integrating the left-hand side of the above inequality w.r.t. $\gamma$ leads to the contradiction.) Since the left-hand side must sum to unity when integrating w.r.t. $\nu$, it further follows that
$\int\frac{e^{u(x,\theta)}\left(\frac{d\nu(x)}{d\phi(x)}\right)^{-\alpha}}{\int e^{u(\xi,\theta)}\left(\frac{d\nu(\xi)}{d\phi(\xi)}\right)^{-\alpha}\>d\nu(\xi)}-1\>d\mu=0$
for a.e. $x$. Then, $a(\xi)=0$ and $b(\theta)=\log\left(\int e^{u(\xi,\theta)}\>d\phi(\xi)^{\alpha}\>d\nu(\xi)^{1-\alpha}\right)$ satisfy the Schrödinger equations for the marginal $\nu$. Thus, it follows if $\nu$ maximizes $f$,
$dP(\xi,\theta)=\frac{e^{u(\xi,\theta)}d\phi(\xi)^{\alpha}d\nu(\xi)^{1-\alpha}}{e^{b(\theta)}}$
is not only a valid probability measure, but also $P|_{\xi}=\nu$, $P|_{\theta}=\mu$, and $P$ maximizes $U$ subject to those constraints. ∎
Proof of corollary˜15.1.
If $P,P^{\prime}$ both maximize $U$, then $P|_{\xi},P^{\prime}|_{\xi}$ both maximize $f$. Since $f$ is strictly concave, it follows that the marginals are equal: $P|_{\xi}=P^{\prime}|_{\xi}=\nu$. Because the joint must satisfy the MNL formula, it follows $P$ and $P^{\prime}$ are just versions of one another. ∎
The Envelope
As suggested, the Envelope Theorem provides a nice heuristic interpretation for $a_{\nu}(x)$. It is left to show, however, that $V(\nu_{\varepsilon})$ or $V(\nu_{\varepsilon})$ can be written as the supremum of a function whose derivative is equicontinuous [5].
If $b_{\nu}$ is a Schrödinger potential, then plugging in $a_{\nu}=\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b_{\nu}}\>d\mu\right)$ gives
$V(\nu)=\int\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b_{\nu}}\>d\mu\right)\ d\nu-\int b_{\nu}\>d\mu$
Normalize $u$ such that $\underline{u}\geq 0$. We are allowed to perform one normalization on $b$, so assume $\int b_{\nu}\>d\mu=0$. Then,
$V(\nu)=-\alpha D_{KL}(\nu\|\phi)+\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)\>d\nu$
By construction of $V(\nu)=\sup_{P\in{\cal Q}(\nu)}U(P)$,
$\overline{u}-\alpha D_{KL}(\nu\|\phi)\geq V(\nu)$
so that
$\overline{u}\geq\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)\>d\nu$
Because $\underline{u}>0$, we have
$\log\left(\int e^{-b_{\nu}}\>d\mu\right)\leq\log\left(\int e^{\underline{u}-b_{\nu}}d\mu\right)\leq\int\log\left(\int e^{u-b_{\nu}}\>d\mu\right)d\nu\leq\overline{u}$
Define $B$ as
$B=\bigg{\{}b\in L^{1}(\mu):\int b\>d\mu=0,\ \log\left(\int e^{-b}\;d\mu\right)\leq\overline{u}\bigg{\}}$
The dual can be written as
$-V(\nu)$ $=\sup_{a\in L^{1}(\phi),\ b\in L^{1}(\mu)}\int-a\>d\mu+\int-b\>d\nu-\iint e^{u(\xi,\theta)-a-b}\>d\phi^{\alpha}d\nu^{1-\alpha}\>d\mu+1$
$=\sup_{b\in B}\int-\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b}\>d\mu\right)\ d\nu$
$=\alpha D_{KL}(\nu\|\phi)+\sup_{b\in B}\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu$
since $B$ imposes no binding restrictions.
Define $\nu_{\varepsilon}=(1-\varepsilon)\nu+\varepsilon\gamma$ such that $d\nu_{\varepsilon}/d\phi$ is bounded uniformly away from 0 and $\infty$ on $\varepsilon\in(-r,r)$. Define $g:B\times(-r,r)\to\mathbb{R}$ by
$g(b,\varepsilon)=\alpha D_{KL}(\nu\|\phi)+\int-\log\left(\int e^{u-b}\>d\mu\right)\ d\nu$
Then,
$-V(\nu_{\varepsilon})=\sup_{b\in B}g(b,\varepsilon)$
By first-step orthogonality, $\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}D_{KL}(\nu_{\varepsilon}\|\phi)=\int\log\left(\frac{d\nu_{t}}{d\phi}\right)\>d(\gamma-\nu)$, and so
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)=\int\alpha\log\left(\frac{d\nu_{t}}{d\phi}\right)-\log\left(\int e^{u-b}\>d\mu\right)\ d(\gamma-\nu)$
By assumption, $\log(d\nu_{t}/d\phi)$ is uniformly bounded. By Jensen’s inequality,
$\log\left(\int e^{u-b}\>d\mu\right)\geq\int u-b\>d\mu=\underline{u}$
Conversely, since $\int e^{-b}\>d\mu\leq e^{\overline{u}}$, by Holder’s inequality
$\int e^{u-b}\>d\mu\leq e^{\overline{u}}\int e^{-b}\>d\mu=e^{2\overline{u}}$
so $\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)$ is uniformly bounded across $b\in B$.
Furthermore, notice that for any $b$,
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)-\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=s}g(b,\varepsilon)=\int\alpha\log\left(\frac{d\nu_{t}}{d\nu_{s}}\right)\>d(\gamma-\nu)$
i.e. it does not depend on $b$. Thus, $\bigg{\{}\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=t}g(b,\varepsilon)\bigg{\}}_{b\in B}$ is equicontinuous.
Lastly,
$\frac{d}{d\varepsilon}\bigg{|}_{\varepsilon=0}g(b_{\nu},\varepsilon)=\int-\log\left(\left(\frac{d\nu}{d\phi}\right)^{-\alpha}\int e^{u-b_{\nu}}\>d\mu\right)\ d(\gamma-\nu)=\int-a_{\nu}\>d(\gamma-\nu)$
Thus,
$\frac{dV(\nu_{\varepsilon})}{d\varepsilon}\bigg{|}_{\varepsilon=0}=\int a_{\nu}\>d(\gamma-\nu)$
If $X$ is a metric space with full support and $a_{\nu}$ is continuous and bounded, then we can construct a sequence of probabilities $\gamma_{n}$ which converges weakly to the Dirac measure $\delta_{x}$, so that
$\frac{dV(\nu_{\varepsilon,n})}{d\varepsilon}\bigg{|}_{\varepsilon=0}\to_{n}\int a_{\nu}\>d(\delta_{x}-\nu)=a_{\nu}(x)-\int a_{\nu}\>d\nu$
Sinkhorn’s algorithm
In the discrete case, the problem of finding Schrödinger potentials such that eq.˜5 holds is equivalent to the matrix scaling problem. Formally, we can re-write eq.˜5 in the discrete case as
$P(\xi,\theta)=e^{-a_{\nu}(\xi)}\left[e^{u(\xi,\theta)}\phi(\xi)^{\alpha}\nu(\xi)^{1-\alpha}\mu(\theta)\right]e^{-b_{\nu}(\theta)}$
If we think of the middle term as a $|X|\times|\Theta|$ matrix $A$, the Schrödinger potentials as diagonal matrices, and $P$ as a $|X|\times|\Theta|$ matrix such that the marginals are $\nu$ and $\mu$ respectively, we can re-formulate the problem as follows: given a matrix A and a pair of probability vectors $\boldsymbol{\nu}$, $\boldsymbol{\mu}$, find diagonal matrices $\textbf{D}_{1},\textbf{D}_{2}$ such that for $\textbf{P}:=\textbf{D}_{1}\textbf{A}\textbf{D}_{2}$ is a probability matrix whose columns sum to $\sum_{\xi}P_{\xi\theta}=\mu_{\theta}$ and rows sum to $\sum_{\theta}P_{\xi\theta}=\nu_{\xi}$.
Theorem 16 (Sinkhorn’s algorithm).
Set $a_{t}=0$. For $t\geq 0$, set
$b_{t}(\theta)=\log\left(\int e^{u(\xi,\theta)-a_{t}(\xi)}\>d\phi^{\alpha}\>d\nu^{1-\alpha}\right)$
$a_{t+1}(\theta)=\log\left(\int e^{u(\xi,\theta)-b_{t}(\theta)}\>d\mu\right)$
We may define $P_{2t}$ by via the potentials $a_{t}$, $b_{t}$, and $P_{2t-1}$ via the potentials $a_{t},b_{t-1}$.
Then:
  1. 1.
    $a_{t}$ and $b_{t}$ converge pointwise and in their respective $L^{p}$-norms, $1\leq p<\infty$
  2. 2.
    $D_{KL}(P\|P_{t})\to 0$
  3. 3.
    $D_{KL}(P_{t}\|R)\to D_{KL}(P\|R)$
  4. 4.
    $P_{t}\to P$ in total variation
Proof.
By assumption, $u$ and $\log\frac{d\nu}{d\phi}$ are totally bounded. This then follows from [6, Theorem 6.15]. ∎